can someone please help me?
solve each system by the substitution method.
8. 2x+y-x=-3y+5
y+2=x-5

Answer by rfer(16322)

x=-4y+5
y=x-7
-------------
x=-4(x-7)+5
x=-4x+28+5
5x=33
x=6.6
-----------
6.6=-4y+5
1.6=-4y
y=1.6/-4
y=-0.4
--------------
-0.4=6.6-7
(6.6,0.4)